OpenStudy (anonymous):

Find the 31st term of the following sequence. 9, 15, 21, ... 186 189 195

OpenStudy (raden):

use formula : an = a1+(n-1)d

OpenStudy (raden):

from information above, a1=9, d=a2-a1=a3-a2=15-9=21-15=6 so, the 31st term is a31 = a1 + (31-1)6 a31=9 + 30(6) a31 = ....?

OpenStudy (raden):

got it ?

OpenStudy (raden):

what ur answer @tlaquita